Cullumber, Inc., sells two types of water pitchers, plastic and glass. Plastic pitchers cost the company $35 and are sold for $50. Glass pitchers cost $44 and are sold for $65. All other costs are fixed at $2,958,228 per year. Current sales plans call for 42,140 plastic pitchers and 126,420 glass pitchers to be sold in the coming year.

How many pitchers of each type must be sold to break even in the coming year? (Use contribution margin per unit to calculate breakeven units.)

Glass $

Plastic $

Cullumber, Inc., has just received a sales catalog from a new supplier that is offering plastic pitchers for $33. What would be the new contribution margin per unit if managers switched to the new supplier?

Plastic Glass
Contribution Margin Per Unit

What would be the new breakeven point if managers switched to the new supplier? (Use contribution margin per unit to calculate breakeven units. Round answers to 0 decimal places, e.g. 25,000.)

Plastic Glass
Breakeven in Units
Answer #1

Requirement 1

Units to breakeven

Plastic                  37,926
Glass                113,778

Working

Working Plastic Glass
A Price $                50.00 $          65.00
B Variable Cost per unit $                35.00 $          44.00
C = A - B Contribution Margin $                15.00 $          21.00
D Product Mix 25% 75%
E = C x D Weighted Average Contribution Margin $                   3.75 $          15.75

.

A Total Fixed Cost plus profit $ 2,958,228.00
B Weighted Average Contribution Margin (3.75+15.75) $                19.50
C = A/B Multi Product Break Even point                151,704
C x 25% Plastic                  37,926
C x 75% Glass                113,778

Requirement 2

Plastic Glass
Contribution Margin $                17.00 $          21.00

Working

Working Plastic Glass
A Price $                50.00 $          65.00
B Variable Cost per unit $                33.00 $          44.00
C = A - B Contribution Margin $                17.00 $          21.00

Requirement 3

New breakeven

Plastic                  36,978
Glass                110,934

Working

Working Plastic Glass
A Price $                50.00 $          65.00
B Variable Cost per unit $                33.00 $          44.00
C = A - B Contribution Margin $                17.00 $          21.00
D Product Mix 25% 75%
E = C x D Weighted Average Contribution Margin $                   4.25 $          15.75

.

A Total Fixed Cost plus profit $ 2,958,228.00
B Weighted Average Contribution Margin (4.25+15.75) $                20.00
C = A/B Multi Product Break Even point                147,911
C x 25% Plastic                  36,978
C x 75% Glass                110,934